Dear colleagues,


Here is a summary of today's activites at CSBF.

  • On Sunday early morning around 6:30 AM (CDT), there was a kind of power stoppage for about guesstimated 10-15 minutes and the muon run was stopped. The SFC was rebooted and cdaq_serv program was started but not the instrument power and HV which requires commanding. The data were copied to cdps3.
  • MXR box cabling and labeling were done.
  • Muon run in flight mode started around 11:30 PM.
  • This early morning around 6:30 AM again, the same problem happened and muon run stopped.
  • Asked Chris Field (CSBF correspondent) to check if there were any activities related with those accidents. But the question is why only CREAM instrument power was off at the almost same time within 10 minutes for two days. The CDM machines and CREAM Client PCs did not show any signs of power off. Chris so far hasn't found any related activities at CSBF. Only one difference between last two nights and previous nights are the change of the power outlet which was used for the UPS of the CREAM Power GSE. Prashant still investigating by looking at the log file of the SFC1 which was used during the runs.
  • MXR box shipped back to PSU for a repair and a replacement of the pxr box which gives time out errors from time to time (~ 4 per 200 events).
  • The remote commanding and data flow test with (CSBFWFFUMD SOC) connection started from the morning to the end of the day.
    • At the beginning, there was a problem in the data flow from the SFC1 to the Client. Turned out that the SFC1 setting was not fully updated yet. Switched to the SFC2, the data flow started. Prashant is going to update the SFC1 setting.
    • At UMD side, cdps1 and could not receive the data for about 3 hours due to the firewall setting related with the allowed IP addresses. Even though it is safe to block the non-local network machines, I think that it is better not to block those IP addresses which were already in the setting. This is a lesson learned during today's test.
    • In summary, the data flow part could be tested at the SOC. The remote commanding and Iridium mode tests are to be done in tomorrow which will start from 9 AM (CDT).
  • The CD threshold control cable's BERG connector was repaired by the CSBF person and re-installed.
  • S3 threshold tuning has started. Tried to reduce the threshold values in a few possible channels the script but did not work. Nick suggested new values in different channels. The first ones of each PMTs would be changed from 100 to 150 to reduce the S3Lo rate. Haven't had time to take data due to other priority.
  • Debugging of the reduced hsk data continued by Prashant by using a Iridium connection.
  • SangEun continue working on the SIP interface coding. A problem with the number "3" character being broken in a serial port communication hold his progress. Has asked may experts but no solution yet. Since SIP is the serial port communication, it is the most critical for him to understand and master the serial connection.

Here are things to do.

  • Continuation and completion of the remote commanding and data flow via Iridium mode at UMD SOC via WFF.
  • Power failure investigation. Until the investigation is over, stop the overnight muon run for the safety.
  • S3 threshold value tuning.
  • Reduced hsk debugging.
  • Chercam HSK problem debugging.
  • TCD 12 V HSK attenuation circuit installation.
  • Update of manuals including CPT.
  • Copying and backup of newly taken integration data into cdps3 at UMD and client PCs at CSBF.
  • Adding another Ethernet card to cdps5 for a dual connection to local network and to outside.

Please let us know if you have any questions/comments/suggestions.

Cheers, MooHyun.


MooHyun, July 21, 2008

UMD